OpenStudy (anonymous):

Find all values of k so that each polynomial can be factored using integers. 1.) x^2+kx-19 2.) x^2 - 8x+k, k>0

OpenStudy (klimenkov):

what means using integers?

OpenStudy (klimenkov):

Integer roots or integer value of k?

OpenStudy (anonymous):

I don't know...I just wrote word for word for what it says in the book...All I know is that for #1 the answers will be in integers...

OpenStudy (klimenkov):

For the first one: -18, 18.

OpenStudy (anonymous):

How do you solve it? Distributive...factor?

OpenStudy (klimenkov):

Do you know Vieta's Theorem?

OpenStudy (anonymous):

Not yet, The questions I got was from the book on the chapter of Quadratic equations, and it said it required Higher-order thinking skills...

OpenStudy (anonymous):

So you factor it out into 2 binomials?

OpenStudy (klimenkov):

Vieta's Theorem is not so hard to get. I advice you to get it and try to use for your task.

OpenStudy (anonymous):

I'm not so sure my teacher wants me to apply Vieta's Theorem yet, because she hates it when we move on ahead of her teachings...btw she can't teach =_=

OpenStudy (anonymous):

Well I have to leave for school in a little while, there's a chemical equation quiz waiting for me...

OpenStudy (anonymous):

Thanks for helping me~

OpenStudy (anonymous):

Hehe why is that? Well I'm sorry I have to go to school now..

OpenStudy (anonymous):

See you guys around^^ Bye

OpenStudy (klimenkov):

If you want to factor \(x^2+kx-19\), I hope you will think it will be \((x-a)(x-b)\). So, when expanding: \(x^2-(a+b)x+ab\). Now just look at the coefficients. \(k=-(a+b), -19=ab\). You want only integers. From the second equality it will be only \(a=1,b=-19\) or \(a=19,b=-1\), because 19 is a prime number. Now try to find \(k\) in both cases.

OpenStudy (klimenkov):

Just try to explain. Not so sure he got it.

OpenStudy (klimenkov):

Or she.

OpenStudy (klimenkov):

If \(-(a+b)=-8\), so \(a+b=8\). And \(ab=k\), we get \(a(8-a)=k>0\). It will be if \(-a^2+8a>0\). Answer: 1, 2, 3, 4, 5, 6, 7.